OpenStudy (anonymous):

compute the derivative y=ln((x+1)^2(3x-1)^3)

OpenStudy (anonymous):

\[y=\ln((x+1)^2(3x-1)^3)\]

OpenStudy (slaaibak):

\[{1 \over (x+1)^2 (3x-1)^3} \times [2\times(x+1)(3x-1)^3 + 3 \times 3 \times (3x-1)^2(x+1)^2] \]

OpenStudy (anonymous):

ok I see

OpenStudy (slaaibak):

or easier: split it up into the following \[2\ln(x+1) + 3\ln(3x-1)\] then derive: \[{2 \over {x+1}} +{ 9 \over 3x-1}\]

OpenStudy (anonymous):

so would that be the answer

OpenStudy (slaaibak):

yes

OpenStudy (anonymous):

do you know how to do \[y=x \sqrt{6}\]

OpenStudy (slaaibak):

yes

OpenStudy (anonymous):

can you show me how to do it? if you have time.

OpenStudy (slaaibak):

well, since \[\sqrt6\] is a constant, the answer would be \[\sqrt6\]

OpenStudy (anonymous):

thanks @slaaibak